Work out the current for a television with the power of 0.8 kW if it uses 200 and 3V mains supply give your answer to one decimal place.
1
Expert's answer
2019-12-05T10:27:08-0500

The power of electrical current is given by


"P=VI"

So, the electrical current


"I=\frac{P}{V}"

"I_1=\frac{800\:\rm W}{200\:\rm V}=4.0\:\rm A"

"I_2=\frac{800\:\rm W}{3\:\rm V}=266.7\:\rm A"